# Long-range Meta-path Search on Large-scale Heterogeneous Graphs

### v4

@@ 7 Conclusion
This work presented Long-range Meta-path Search through Progressive Sampling (LMSPS) to tackle the challenges of leveraging long-range dependencies in large-scale heterogeneous graphs, i.e., reducing computational costs while effectively utilizing information and addressing the over-smoothing issue. Based on our two observations, i.e., a few meta-paths dominate the performance, and certain meta-paths can have negative impact on performance, LMSPS introduced a progressive sampling search algorithm and a sampling evaluation strategy to automatically identify effective meta-paths, thus reducing the exponentially growing number of meta-paths to a manageable constant. Extensive experiments demonstrated the superiority of LMSPS over existing methods, particularly on sparse heterogeneous graphs that require long-range dependencies. By employing simple MLPs and complex meta-paths, LMSPS offers a novel direction that emphasizes data-dependent semantic relationships rather than relying solely on sophisticated neural architectures. The reproducibility and limitations are discussed in Appendix E and Appendix F, respectively.
Appendix
@@
In the Appendix, we provide additional details and results not included in the main text due to space limitations. We first show the training objective and search algorithm of LMSPS and explore the algorithm’s effectiveness. Secondly, for interpretability, the searched meta-paths are listed and analyzed. Thirdly, we display more experimental details, explore the training efficiency of LMSPS and compare LMSPS with baselines on Freebase. Ablation studies and hyperparameter studies are also conducted. Finally, we discuss the reproducibility and limitations of our research work.
Appendix A Searching and Training Objective
@@
Following the convention (Lv et al., 2021; Li et al., 2023; Yang et al., 2023; Mao et al., 2023), we concentrate on semi-supervised node classification under the transductive setting and leave other downstream tasks related to heterogeneous graph representation learning as future work. An MLP layer is appended after the main module of LMSPS, which reduces the dimension of node-level representations of a heterogeneous graph to the number of classes:
@@
7:     Calculate the current search space $\mathbb{S}_{C}$ based on Equations 3 and 4
@@
9:     Determine the indexes of sampled meta-paths based on Equation 6
@@
Appendix B The Search Algorithm
Our search stage aims to discover the most effective meta-path set from all target-node-related meta-paths, severely challenging the efficiency of searching. Take ogbn-mag as an example. The number of target-node-related meta-paths is 226 under the maximum hop $6$, and we need to find the most effective meta-path set with size 30. Because different meta-paths could be noisy or redundant to each other, top-30 meta-paths are not necessarily the optimal solution when their importance is calculated independently. Therefore, the total number of possible meta-path sets is $C_{226}^{30}\approx 10^{37}$. Such a large search space is hard to solve efficiently by traditional RL-based algorithms (Zhong et al., 2020; Ning et al., 2022) or evolution-based algorithms (Xie & Yuille, 2017; Real et al., 2017).
To overcome this challenge, our LMSPS first uses a progressive sample algorithm to shrink the search space size from 226 to 60, then utilizes a sampling evaluation strategy to discover the best meta-path set. In each iteration, we only uniformly sample meta-paths $M$ from the whole search space for parameter updates, so the search cost is relevant to $M$, which is a predefined small number, rather than $K$. Because the search stage has many iterations and the initial values of architecture parameters are the same, all architecture parameters will be updated multiple times and the relative importance can be learned during training, making the total search cost similar to training a single HGNN once. Specifically, for ogbn-mag, LMSPS can finish searching in two hours.
Except for efficiency, LMSPS can also overcome the over-squashing issue (Alon & Yahav, 2021) when utilizing long-range dependency. Over-squashing means the distortion of messages being propagated from distant nodes, which has been heuristically attributed to graph bottlenecks where the number of $l$-hop neighbors grows rapidly with $l$. In LMSPS, we set the number of searched meta-paths $M=30$ for all datasets, which is independent of $l$. Because the exponential meta-paths in metapath-based methods correspond to exponential receptive fields in metapath-free methods, the constant $M$ means approximately constant $l$-hop neighbors, i.e., LMSPS selectively aggregate effective neighbors for each target-net. Therefore, the over-squashing issue is overcome.
Appendix C Interpretability of Searched Meta-paths
@@
Table 8: Statistics of datasets used in this paper.
Dataset #Nodes #Node types #Edges #Edge types Target #Classes
DBLP ${}~{}~{}~{}26,128$ $4$ ${}~{}~{}~{}239,566$ $6$ author $4$
IMDB ${}~{}~{}~{}21,420$ $4$ ${}~{}~{}~{}~{}~{}86,642$ $6$ movie $5$
ACM ${}~{}~{}~{}10,942$ $4$ ${}~{}~{}~{}547,872$ $8$ paper $3$
ogbn-mag $1,939,743$ $4$ $21,111,007$ $4$ paper $349$
ogbn-mag-50 $1,939,743$ $4$ ${}~{}9,531,403$ $4$ paper $349$
ogbn-mag-20 $1,939,743$ $4$ ${}~{}7,975,003$ $4$ paper $349$
ogbn-mag-10 $1,939,743$ $4$ 6,610$,46$ 610 , 464 $4$ paper $349$
ogbn-mag-5 $1,939,743$ $4$ ${}~{}4,958,941$ $4$ paper $349$
@@
Appendix D More Experiments
D.1 Dataset Details
We evaluate our method on the famous large-scale dataset ogbn-mag from OGB challenge (Hu et al., 2021), and three widely-used heterogeneous graphs including DBLP, IMDB, and ACM from HGB benchmark (Lv et al., 2021). For the ogbn-mag dataset, we use the official data partition, where papers published before 2018, in 2018, and since 2019 are nodes for training, validation, and testing, respectively. The datasets from HGB follow a transductive setting, where all edges are available during training, and target type nodes are divided into $24\%$ for training, $6\%$ for validation, and $70\%$ for testing. In addition, we construct four sparse datasets to evaluate the performance of LMSPS by reducing the maximum in-degree related to edge type in ogbn-mag. The four sparse datasets have the same number of nodes with ogbn-mag. Follow the convention (Lv et al., 2021; Yang et al., 2023; Mao et al., 2023), we show the details of all datasets in Table 8.
D.2 More Training Details
We use Pytorch (Paszke et al., 2019) to run all experiments on one Tesla V100 GPU with 16GB GPU memory. A two-layer MLP is adopted for each meta-path in the feature projection step and the hidden size is $512$. All network weights are initialized by the Xavier uniform distribution (Glorot & Bengio, 2010) and are optimized with Adam (Kingma & Ba, 2015) during training. In the search stage, $\lambda$ is $1$ during the first $20$ epochs for warmup and decreases to 00 linearly. $\tau$ linearly decays with the number of epochs from $8$ to $4$. The learning rate is $0.001$ for all search stages and HGB training stage, and $0.003$ for ogbn-mag training stage. The weight decay is always 00. For the initial search space, we simply preset the maximum hop and use all target-node-related meta-paths no more than this maximum hop.
As we adopt the results of most baselines published in OGB leaderboard (Hu et al., 2021) and HGB (Lv et al., 2021), we only need to run the baselines whose results are not reported in OGB leaderboard and HGB. For these methods, we employ their original settings. When involving hyperparameter analysis like in Figure 3, we only change the corresponding hyperparameters for a fair comparison.
Method DBLP IMDB ACM
HAN $95.44\pm 0.14$ $65.95\pm 0.31$ $90.66\pm 0.30$
GTN $95.33\pm 0.05$ $65.99\pm 0.16$ $90.66\pm 0.30$
DARTS $95.35\pm 0.17$ $66.23\pm 0.14$ $93.45\pm 0.13$
SPOS $95.41\pm 0.43$ $67.10\pm 0.29$ $93.64\pm 0.37$
DiffMG $95.45\pm 0.49$ $66.98\pm 0.37$ $93.61\pm 0.45$
PMMM $95.48\pm 0.27$ $67.49\pm 0.24$ $93.74\pm 0.22$
LMSPS $95.66\pm 0.20$ $68.70\pm 0.26$ $94.69\pm 0.36$
@@ D.2 More Training Details
D.3 Effectiveness of the Search Algorithm
This subsection explores the effectiveness of the search algorithm. In our architecture, our meta-paths are replaced by those meta-paths discovered by other methods. DARTS (Liu et al., 2019a) is the first differentiable search algorithm in neural networks. SPOS (Guo et al., 2020) is a classic singe-path differentiable algorithm. DARTS and SPOS aim to search operations, like $3\times 3$ convolution, in convolutional neural networks. DiffMG (Ding et al., 2021) and PMMM (Li et al., 2023) search for meta-graphs instead of meta-paths and cannot search on large-scale datasets. We ignore these differences and focus on the algorithms. The derivation strategies of the four methods are unsuitable for discovering multiple meta-paths. So we changed their derivation strategies to ours to improve their performance. We report the Micro-F1 scores of DBLP, IMDB, and ACM in Table 9. It can be seen that LMSPS shows the best performance, verifying the effectiveness of our search algorithm.
D.4 Comparison on Training Efficiency
@@
We show the time cost and parameters of LMSPS and the advanced baselines on DBLP and ACM in Figure 4. Following the convention (Lv et al., 2021; Yang et al., 2023), we measure the average time consumption of one epoch for each model. The area of the circles represents the parameters. The hidden size is set to $512$ and the maximum hop or layer is $6$ for DBLP and $5$ for ACM for all methods to test the training time and parameters under the same setting. Some methods perform quite poorly under the large maximum hop or layer. So we show the performance from Table 3 of the main text, which is the results under their best settings. Figure 4 shows that LMSPS has advantages in both training efficiency and performance. Our searched meta-paths are universal after searching once and can be applied to other meta-path-based HGNNs based on Table 6. Because other meta-path-based HGNNs don’t include the time for discovering manual meta-paths, we also exclude our search time for discovering meta-paths in Figure 4. The search time is shown in Figure 3 (c) of the main text.
[…]
Method Freebase
RSHN (Zhu et al., 2019) $OOM$
HetSANN (Hong et al., 2020) $OOM$
GTN (Yun et al., 2019) $OOM$
HGT (Hu et al., 2020b) $60.51\pm 1.16$
Simple-HGN (Lv et al., 2021) $66.29\pm 0.45$
HINormer (Mao et al., 2023) --
RGCN (Schlichtkrull et al., 2018) $58.33\pm 1.57$
HetGNN (Zhang et al., 2019) $OOM$
HAN (Wang et al., 2019) $54.77\pm 1.40$
MAGNN (Fu et al., 2020) $OOM$
SeHGNN (Yang et al., 2023) $63.41\pm 0.47$
DiffMG (Ding et al., 2021) --
PMMM (Li et al., 2023) --
LMSPS $66.52\pm 0.49$
D.5 Further Comparison on Freebase Dataset
Compared with DBLP, IMDB, and ACM, Freebase is less employed by recent HGNNs due to the appearance of ogbn-mag. We ignore Freebase in the main text because of the space limitation. Here, we compare LMSPS with the related baselines on Freebase from HGB (Lv et al., 2021) in Table 10 using the settings in Sections 6.2 and Appendix D.2. As we can see, LMSPS outperforms other HGNNs on Freebase, which is consistent with the results in Table 3.
Table 11: Experiments on DBLP, IMDB, ACM, and ogbn-mag to analyze the effects of different blocks in LMSPS. PS means progressive sampling strategy, and SE means sampling evaluation strategy. * means replacing the concatenation operation with the transformer module, and $\dagger$ means employing all meta-paths and transformer module.
Method DBLP IMDB ACM ogbn-mag
LMSPS $95.66\pm 0.20$ $68.70\pm 0.26$ $94.69\pm 0.36$ $54.83\pm 0.20$
LMSPS w/o PS $95.00\pm 0.19$ $66.52\pm 0.37$ $93.14\pm 0.41$ $47.66\pm 0.45$
LMSPS w/o SE $95.48\pm 0.24$ $67.13\pm 0.47$ $94.15\pm 0.31$ $52.94\pm 0.34$
LMSPS * $95.57\pm 0.23$ $68.54\pm 0.16$ $94.63\pm 0.28$ $54.72\pm 0.42$
LMSPS ${{}^{\dagger}}$ $95.38\pm 0.21$ $68.58\pm 0.34$ $94.57\pm 0.39$ $OOM$

### v6

@@ 7 Conclusion
This work presented a novel approach, the Long-range Meta-path Search through Progressive Sampling (LMSPS), to tackle the challenges of leveraging long-range dependencies in large-scale heterogeneous graphs, i.e., reducing computational costs while effectively utilizing information and addressing the over-smoothing issue. Based on our two observations, i.e., a few meta-paths dominate the performance, and certain meta-paths can have negative impact on performance, LMSPS introduced a progressive sampling search algorithm and a sampling evaluation strategy to automatically identify effective meta-paths, hence reducing the exponentially growing number of meta-paths to a manageable constant. Extensive experiments demonstrated the superiority of LMSPS over existing methods, particularly on sparse heterogeneous graphs that require long-range dependencies. By employing simple MLPs and complex meta-paths, LMSPS offers a novel direction that emphasizes data-dependent semantic relationships rather than relying solely on sophisticated neural architectures. The reproducibility and limitations are discussed in Appendix F and Appendix G, respectively.
Appendix / Supplemental Material
@@
In the Appendix, we provide additional details and results not included in the main text due to space limitations. We first provide the details of datasets, tasks, baselines, and parameter settings of our experiments. Secondly, we show the search algorithm of LMSPS and analyze the time complexity. Thirdly, for interpretability, the searched meta-paths are listed and analyzed. Fourthly, we explore the training efficiency and search convergence of LMSPS and compare LMSPS with top method combinations on OGBN-MAG leaderboard. Hyperparameter studies are also conducted. Finally, we discuss the reproducibility and limitations of our research work.
Appendix A Dataset and Task
A.1 Dataset Details
We evaluate our method on four widely-used heterogeneous graphs including DBLP, IMDB, ACM, and Freebase from HGB benchmark [32] and the famous large-scale dataset OGBN-MAG from OGB challenge [18]. The datasets from HGB follow a transductive setting, where all edges are available during training, and target type nodes are divided into $24\%$ for training, $6\%$ for validation, and $70\%$ for testing. For the OGBN-MAG dataset, we use the official data partition, where papers published before 2018, in 2018, and since 2019 are nodes for training, validation, and testing, respectively. In addition, we construct four sparse datasets to evaluate the performance of LMSPS by reducing the maximum in-degree related to edge type in OGBN-MAG. The four sparse datasets have the same number of nodes with OGBN-MAG. Follow the convention [32, 51, 34], we show the details of all datasets in Table 7.
A.2 Node Classification Tasks
Following the convention [32, 25, 51, 34], we concentrate on semi-supervised node classification under the transductive setting and leave other downstream tasks related to heterogeneous graph representation learning as future work. Each dataset contains a target node type, and our task is to learn to predict the labels of vertices of this target node type. All the datasets provide fixed data splitting (into training, validation, and test sets) for node classification tasks.
Table 7: Statistics of the datasets.
Dataset #Nodes #Node types #Edges #Edge types Target #Classes
DBLP ${}~{}~{}~{}26,128$ $4$ ${}~{}~{}~{}239,566$ $6$ author $4$
IMDB ${}~{}~{}~{}21,420$ $4$ ${}~{}~{}~{}~{}~{}86,642$ $6$ movie $5$
ACM ${}~{}~{}~{}10,942$ $4$ ${}~{}~{}~{}547,872$ $8$ paper $3$
Freebase ${}~{}~{}180,098$ $8$ ${}~{}~{}~{}1,057,688$ $7$
OGBN-MAG $1,939,743$ $4$ $21,111,007$ $4$ paper $349$
OGBN-MAG-50 $1,939,743$ $4$ ${}~{}9,531,403$ $4$ paper $349$
OGBN-MAG-20 $1,939,743$ $4$ ${}~{}7,975,003$ $4$ paper $349$
OGBN-MAG-10 $1,939,743$ $4$ 6,610$,46$ 610 , 464 $4$ paper $349$
OGBN-MAG-5 $1,939,743$ $4$ ${}~{}4,958,941$ $4$ paper $349$
A.3 Searching and Training Objective
@@
An MLP layer is appended after the main module of LMSPS, which reduces the dimension of node-level representations of a heterogeneous graph to the number of classes:
@@
Appendix B Baselines and Parameter Settings
B.1 Baselines
We compare LMSPS with a large number of HGNN baselines, including MLP [17], GraphSAGE [13], RGCN [39], HAN [46], GTN [55], RSHN [62], HetGNN [56], MAGNN [10], HetSANN [16], GCN [24], GAT [44], , Simple-HGN [32], HGT [19], GraphMSE [29], SlotGAT [61], DiffMG [7], PMMM [25], HINormer [34], NARS [54], space4HGNN [58], SeHGNN [51]. GEMS [14] and RMS-HRec [35] are ignored because they are designed for recommendation tasks. RL-HGNN [59] is omitted due to a lack of source code.
Note that most of these baselines encounter out-of-memory (OOM) issues when dealing with large datasets. However, there are exceptions among them, including MLP, GraphSAGE, HGT, RGCN, NARS, and SeHGNN. The results of MLP and GraphSAGE come from the OGBN-MAG leaderboard. MLP [17] is famous for its lightweight and GraphSAGE [13] uses a neighbor sampling strategy, so they can handle large-scale datasets. HGT is designed with a graph sampling strategy, making it well-suited for managing large datasets. As for RGCN, while the original version does present OOM issues with large datasets, it is often used as a baseline for large-scale HGNNs [54]. NARS [54] and SeHGNN [51] are metapath-based and pre-computation-based HGNNs, allowing them to handle large-scale datasets.
It’s worth noting that in the original paper, SeHGNN refers to the full version incorporating label propagation techniques. For clarity and consistency in our study, we have adjusted the terminology: we refer to the version without label propagation as SeHGNN and the version with label propagation as SeHGNN+LP. Note that label propagation is a general technique [63, 60, 20] that is compatible with most GNNs and is not used in the results of LMSPS and all other baselines in Table 1. Therefore, it is considered only in Appendix E.2, where we demonstrate the successful integration of our model with other techniques.
B.2 Training Settings
For training, we follow the settings of the HGB benchmark [32] and utilize the performance improvement on the validation set as a guide to determine whether the model has been improved. Specifically, if we observe a performance boost on the validation set during training, we update the final model parameters accordingly. Additionally, we adopt the early stopping strategy employed by the HGB benchmark: if no performance improvement on the validation set is observed within a specific count (referred to as patience), the training will be stopped early before reaching the maximum epochs. This approach helps prevent overfitting while also enhancing the computational efficiency of our model training. Following the setting of HGB [32] and OGB [18], we train our model 5 times for HGB and 10 times for OGB with different random seeds and report the mean performance and standard deviation, respectively. We use Pytorch [36] to run all experiments on one Tesla V100 GPU with 16GB GPU memory.
B.3 Parameter Settings for Baselines
For models such as RGCN, HAN, GTN, RSHN, HetGNN, MAGNN, HetSANN, HGT, GCN, GAT, Simple-HGN, and SeHGNN, the literature  [32, 51] provides tuned parameter settings along with corresponding performance on small datasets, namely DBLP, IMDB, ACM, and Freebase. Therefore, we utilize these tuned parameter settings for these models on small datasets and utilize the reported performance.
Conversely, in cases where official results are unavailable (such as NARS on small datasets), or where the experimental settings differ, we employ the official or benchmark implementations  [^1] https://github.com/THUDM/HGB  [^2] https://github.com/dmlc/dgl/tree/master/examples/pytorch/ogb/ogbn-mag  [^3] https://github.com/UCLA-DM/pyHGT  [^4] https://github.com/facebookresearch/NARS  [^5] https://github.com/LARS-research/DiffMG  [^6] https://github.com/JHL-HUST/PMMM  [^7] https://github.com/ICT-GIMLab/SeHGNN  [^8] https://github.com/Ffffffffire/HINormer  [^9] https://github.com/scottjiao/SlotGAT_ICML23 of the baseline models. We meticulously fine-tune their hyperparameters to the best of our capabilities. In instances involving hyperparameter analysis, such as in Figure 3, we only modify the relevant hyperparameters to ensure a fair comparison.
B.4 Parameter Settings for LMSPS
We set the number of selected meta-paths $M=30$ for all datasets. The final search space $V=60$. The maximum hop is $6$ for ogbn-mag, DBLP, $5$ for IMDB, ACM, and $3$ for Freebase. All $K$ architecture parameters $\alpha_{1},\alpha_{2},\cdots,\alpha_{K}$ are initialized as $1$s. For searching in the super-net, we train for $200$ epochs. To train the target-net, we use an early stop mechanism based on the validation performance to promise full training. A two-layer MLP is adopted for each meta-path in the feature projection step, and the hidden size is $512$. All network weights are initialized by the Xavier uniform distribution [11] and are optimized with Adam [23] during training. In the search stage, $\lambda$ is $1$ during the first $20$ epochs for warmup and decreases to 00 linearly. $\tau$ linearly decays with the number of epochs from $8$ to $4$. The learning rate is $0.001$ for all search stages and HGB training stage, and $0.003$ for OGBN-MAG training stage. The weight decay is always 00. For the initial search space, we simply preset the maximum hop and use all target-node-related meta-paths no more than this maximum hop.
B.5 Evaluation Metrics
To assess the performance of the models, we employ evaluation metrics consistent with those employed in baseline models. The metrics are chosen as follows. For small datasets, DBLP, IMDB, ACM, and Freebase, we adhere to the evaluation standards established by the HGB benchmark [32]. The metrics reported for these datasets are Macro-F1 and Micro-F1 scores, which evaluate the classification performance. For the OGBN-MAG dataset, evaluation follows the methodology outlined in NARS [54] and SeHGNN [51], where the classification accuracy score is reported for this dataset.
@@ B.5 Evaluation Metrics
7:     Calculate the current search space $\mathbb{S}_{C}$ based on Equations 3
@@
9:     Determine the set of indexes of sampled meta-paths $\mathbb{S}$ based on Equation 4
@@
Appendix C Algorithm
C.1 The Search Algorithm
Our search stage aims to discover the most effective meta-path set from all target-node-related meta-paths, severely challenging the efficiency of searching. Take OGBN-MAG as an example. The number of target-node-related meta-paths $K$ is 226 under the maximum hop $6$, and we need to find the most effective meta-path set with size 30. Because different meta-paths could be noisy or redundant to each other, top-30 meta-paths are not necessarily the optimal solution when their importance is calculated independently. Therefore, the total number of possible meta-path sets is $C_{226}^{30}\approx 10^{37}$. Such a large search space is hard to solve efficiently by traditional RL-based algorithms [59, 35] or evolution-based algorithms [48, 37].
To overcome this challenge, our LMSPS first uses a progressive sample algorithm to shrink the search space size from 226 to 60, then utilizes a sampling evaluation strategy to discover the best meta-path set with the lowest validation loss, which is more effective than architecture parameters [45]. In each iteration, we only uniformly sample meta-paths $M$ from the whole search space for parameter updates, so the search cost is relevant to $M$, which is a predefined small number, rather than $K$. Because the search stage has many iterations and the initial values of architecture parameters are the same, all architecture parameters will be updated multiple times and the relative importance can be learned during training, making the total search cost similar to training a single HGNN once. Specifically, for OGBN-MAG, LMSPS can finish searching in two hours.
Except for efficiency, LMSPS can also overcome the over-squashing issue [1] when utilizing long-range dependency. Over-squashing means the distortion of messages being propagated from distant nodes, which has been heuristically attributed to graph bottlenecks where the number of $l$-hop neighbors grows rapidly with $l$. In LMSPS, we set the number of searched meta-paths $M=30$ for all datasets, which is independent of $l$. Because the exponential meta-paths in metapath-based methods correspond to exponential receptive fields in metapath-free methods, the constant $M$ means approximately constant $l$-hop neighbors, i.e., LMSPS selectively aggregate effective neighbors for each target-net. Therefore, the over-squashing issue is overcome.
HAN $O(N(rd)^{l}F^{2})$ $O(N(rd)^{l}F)$ $O(Nr^{l}F^{2})$ $O(N(rd)^{l}F^{2})$
Simple-HGN $O(N(rd)^{l}F^{2})$ $O(N(rd)^{l}F)$ $O(N(rd)^{l}F^{2})$
Simple-HGN $O(NrdlF^{2})$ $O(NrdlF)$ $O(NrdlF^{2})$
SeHGNN $O(Nr^{l}F^{2})$ $O(Nr^{l}F^{2}+Nr^{2l}F)$ $O(N(r^{l}F^{2}+r^{2l}F))$
LMSPS-search $O(NMF^{2})$ $O(NMF^{2})$ $O(NMF^{2})$
LMSPS-train $O(NMF^{2})$ $O(NMF^{2})$ $O(NMF^{2})$
C.2 Time Complexity Analysis
Following the convention [5, 51], we compare the time complexity of LMSPS with HAN [46], Simple-HGN [32], and SeHGNN [51] under mini-batch training with the total $N$ target nodes. All methods employ $l$-hop neighborhood. For simplicity, we assume that the number of features is fixed to $F$ for all layers. The average degree of each node is $rd$, where $r$ is the number of edge types and $d$ is the number of edges connected to the node for each edge type. The complexity analysis is summarized in Table 8. Because HAN and Simple-HGN require neighbor aggregation during training, and the number of neighbors grows exponentially with hops. So, they have neighbor aggregation costs of $O((rd)^{l})$. SeHGNN employs a pre-processing step to avoid the training cost of neighbor aggregation. However, the exponential meta-paths cause SeHGNN to suffer from $O(r^{l})$ costs in semantic aggregation. Unlike the above methods, LMSPS samples $M$ meta-paths in each iteration of the search stage and employs $M$ effective meta-paths in the training stage to avoid exponential costs. Generally, we have $O((rd)^{l})\gg O(r^{l})=O(K)\gg O(M)$ when maximum hop $l$ is large. The time complexity of LMSPS is a constant when $N$ and $F$ are determined, which is the key point for utilizing long-range meta-paths.
@@ C.2 Time Complexity Analysis
Appendix D Interpretability of Searched Meta-paths
@@
Appendix E More Experiments
@@
E.1 Comparison on Training Efficiency
@@
Following the convention [32, 51], we show the time cost and parameters of LMSPS and the advanced baselines on DBLP and ACM in Figure 4. We measure the average time consumption of one epoch for each model. The area of the circles represents the parameters. The hidden size is set to $512$ and the maximum hop or layer is $6$ for DBLP and $5$ for ACM for all methods to test the training time and parameters under the same setting. Some methods perform quite poorly under the large maximum hop or layer. So we show the performance from Table 1 of the main text, which is the results under their best settings. Figure 4 shows that LMSPS has advantages in both training efficiency and performance. Our searched meta-paths are universal after searching once and can be applied to other metapath-based HGNNs based on Table 5. Because other metapath-based HGNNs don’t include the time for discovering manual meta-paths, we also exclude our search time for discovering meta-paths in Figure 4. The search time is shown in Figure 3 (c) of the main text.
Table 10: Performance of top mothod combinations on OGBN-MAG leaderboard.
Method OGBN-MAG
Validation Accuracy Test Accuracy
SAGN [40]+LP $52.25\pm 0.30$ $51.17\pm 0.32$
GAMLP [57]+LP $53.23\pm 0.23$ $51.63\pm 0.22$
SeHGNN+LP $55.95\pm 0.11$ $53.99\pm 0.18$
LMSPS +LP $\textbf{56.98}\pm\textbf{0.10}$ $\textbf{55.10}\pm\textbf{ 0.11}$
SeHGNN+LP+MS $58.70\pm 0.08$ $56.71\pm 0.14$
SeHGNN+LP+MS+Emb $59.17\pm 0.09$ $57.19\pm 0.12$
LMSPS +LP+MS $\textbf{59.51}\pm\textbf{0.07}$ $\textbf{57.84}\pm\textbf{0.22}$
E.2 Comparison with Top Method Combinations
The OGBN-MAG benchmark dataset is associated with a public leaderboard. In Table 10, we compare our method combination against top-performing approaches (method combinations) on the leaderboard. These methods integrate several general techniques, such as Label Propagation (LP), Multi-Stage Training (MS), and pre-trained Embeddings (Emb). By integrating LP and MS, LMSPS+LP+MS outperforms SeHGNN+LP+MS by a large margin. SeHGNN shows the results with extra embeddings as an enhancement. Though we cannot conduct a fair comparison with this trick due to the untouchability of its embedding file, LMSPS still outperforms their best result.
Table 11: The performance of LMSPS under different searching epochs.
Dataset $1$ $5$ $10$ $20$ $40$ $60$ $80$ $100$ $120$ $140$ $160$ $180$ $200$
DBLP $38.32$ $68.75$ $71.20$ $72.03$ $71.05$ $70.56$ $73.12$ $73.70$ $73.56$ $73.85$ $73.84$ $74.36$ $74.04$
IMDB $43.48$ $49.78$ $53.88$ $58.48$ $57.03$ $58.40$ $58.05$ $58.65$ $57.54$ $58.70$ $58.59$ $58.72$ $58.81$
ACM $79.03$ $85.43$ $87.42$ $86.75$ $88.30$ $89.85$ $88.96$ $89.40$ $88.52$ $89.85$ $89.70$ $89.85$ $89.91$
Freebase $18.37$ $29.55$ $39.62$ $39.58$ $44.43$ $46.57$ $47.52$ $48.33$ $54.59$ $55.81$ $55.52$ $55.66$ $55.79$
OGBN-MAG $14.21$ $26.20$ $32.41$ $37.00$ $40.82$ $45.89$ $41.74$ $44.00$ $48.40$ $48.55$ $48.70$ $48.43$ $48.59$
E.3 Convergence of the Search Stage
As the search stage of LMSPS relies on sampling, we conduct experiments to explore the convergence of the search stage. Because the test set is unavailable in the search stage to avoid label leakage, we show the Micro-F1 scores of the validation set of DBLP, IMDB, ACM, Freebase and test accuracy of the validation set of OGBN-MAG under different searching epochs in Table 11. As we can see, LMSPS coverages during $100$ to $180$ epochs for all datasets. The main reason is that the search space size is progressively shrunk from $K$ to $2M$ under the progressive sampling algorithm. Most negative meta-paths have been removed from the search space in the second half of the search stage.